# Unifying 3D Vision-Language Understanding via Promptable Queries

### v1

@@ Abstract
A unified model for 3D vision-language (3D-VL) understanding is expected to take various scene representations and perform a wide range of tasks in a 3D scene. However, a considerable gap exists between existing methods and such a unified model, due to the independent application of representation and insufficient exploration of 3D multi-task training. In this paper, we introduce PQ3D, a unified model capable of using Promptable Queries to tackle a wide range of 3D-VL tasks, from low-level instance segmentation to high-level reasoning and planning. This is achieved through three key innovations: (1) unifying various 3D scene representations (i.e., voxels, point clouds, multi-view images) into a shared 3D coordinate space by segment-level grouping, (2) an attention-based query decoder for task-specific information retrieval guided by prompts, and (3) universal output heads for different tasks to support multi-task training. Tested across ten diverse 3D-VL datasets, PQ3D demonstrates impressive performance on these tasks, setting new records on most benchmarks. Particularly, PQ3D improves the state-of-the-art on ScanNet200 by 1.8% (AP), ScanRefer by 5.4% (acc@0.5), Multi3DRefer by 11.7% (F1@0.5), and Scan2Cap by 13.4% (CIDEr@0.5). Moreover, PQ3D supports flexible inference with individual or combined forms of available 3D representations, e.g., solely voxel input.
@@ 1 Introduction
Recent advancements in embodied artificial intelligence have emphasized the importance of connecting 3D scene understanding with natural language [22, 34, 52, 79, 32]. This step is crucial for embodied agents to accurately understand and execute human instructions in real-world scenarios [6, 59]. In recent years, numerous tasks and datasets for benchmarking 3D scene understanding with languages have been proposed, including 3D semantic segmentation [60], 3D vision-language (3D-VL) reasoning (referring [9, 3, 74, 2], question answering [4, 49, 76], and captioning [13]), and open-vocabulary 3D understanding [27, 64, 39].
@@
We conduct extensive experiments on ten 3D-VL datasets, including ScanNet200 [60] / Replica [63] for instance segmentation, ScanRefer [9] / ReferIt3D [3] / Multi3DRef [74] for visual grounding, ScanQA [4] / SQA3D for question-answering, Scan2Cap [13] for dense captioning, and ObjNav from CortexBench [50] for embodied navigation. The proposed PQ3D achieves impressive results across these tasks, setting new records on most tasks as shown in Fig. 2(a). For example, our model boosts the state-of-the-art on ScanNet200 by 1.8% (AP), ScanRefer by 5.4% (acc@0.5), Multi3DRef by 11.7% (F1@0.5), and Scan2Cap by 13.4% (CIDEr@0.5). More importantly, PQ3D is the first all-in-one model capable of handling all these tasks simultaneously. The proposed model also shows zero-shot capability with novel prompt types; for instance, we can prompt it with an image sketch to locate all related objects in a scene as in Fig. 1.
Figure 2: Comparison between PQ3D and other models. (a) When comparing PQ3D to other state-of-the-art (SOTA) methods, PQ3D demonstrates superior performance on most tasks. (b) Previous models have been designed for specific tasks and representations, often limiting the potential for developing a unified model. (c) Our PQ3D offers the unique ability to accommodate various input representations in a flexible way, effectively addressing a wide range of tasks.
@@ 3.1 Task Prompt Encoding
There are three widely used representations for 3D scenes: point clouds, voxel grids, and multi-view images, which have their unique advantages in different tasks. For example, voxels offer a uniform, grid-like structure ideal for instance segmentation [60, 62], point clouds provide detailed and accurate spatial information crucial for visual grounding [79, 11], and multi-view images include a rich visual context beneficial for open-vocabulary scene understanding [53, 64]. These representations can be easily acquired from RGB-D streams, the primary input source for reconstructing 3D indoor scenes.
To simultaneously handle various tasks in a 3D scene, we aim to achieve seamless unification of these representations. To accomplish this, we first group the dense 3D points into larger segments through unsupervised graph-based segmentation [24]. The grouped segments are much fewer than the original points, making the training feasible by reducing the number of tokens in cross-attention. Then, we encode these scene representations by the corresponding encoders and pool the encoded features to the grouped segments in total of $M$.
@@ 3.3 Prompt-guided Query Learning
Following [15, 62], we adopt masked attention when cross-attending to the scene features, which restricts the attention to localized features centered around the query for faster convergence and improved performance. To support flexible inference when only some representations are available, we randomly drop out some scene features in masked-attention computation during training. After $N$ decoder layers, we expect the final instance queries $\mathbf{Q}$ to have collected enough information for solving the given task.
@@ 4.1 Experimental setting
Training Details The training procedure consists of two stages. In the first stage, we train the model with instance segmentation alone on ScanNet200 for 600 epochs. At this stage, instance segmentation is trained with a classification head on instance queries, instead of prompting object categories. At the second stage, we load the checkpoint from stage 1 and train the model on the full training set using the training objectives in Eq. 7 for 50 epochs. We set the hidden dimension $D$ to 768, and query decoder layer $N$ to 4. We utilize the AdamW optimizer with a learning rate of 1e-4, batch size of 16, $\beta_{1}$ = 0.9, and $\beta_{2}$ = 0.98. The loss balance weights $\lambda_{\mathrm{mask}},\lambda_{\mathrm{gen}}$ are set to 1, and $\lambda_{\mathrm{grd}}$ is set to 10. To further demonstrate the capability of PQ3D, we also transfer it to an embodied agent for object navigation using the ObjNav task from CortexBench [50] and instruction-tune it with a large language model (LLM) Vicuna-7B using the instruction-following dataset from [32]. The whole training process is conducted on four NVIDIA A100 GPUs. More details can be found in the appendix.

### v2

@@ Abstract
A unified model for 3D vision-language (3D-VL) understanding is expected to take various scene representations and perform a wide range of tasks in a 3D scene. However, a considerable gap exists between existing methods and such a unified model, due to the independent application of representation and insufficient exploration of 3D multi-task training. In this paper, we introduce PQ3D, a unified model capable of using Promptable Queries to tackle a wide range of 3D-VL tasks, from low-level instance segmentation to high-level reasoning and planning. This is achieved through three key innovations: (1) unifying various 3D scene representations (i.e., voxels, point clouds, multi-view images) into a shared 3D coordinate space by segment-level grouping, (2) an attention-based query decoder for task-specific information retrieval guided by prompts, and (3) universal output heads for different tasks to support multi-task training. Tested across ten diverse 3D-VL datasets, PQ3D demonstrates impressive performance on these tasks, setting new records on most benchmarks. Particularly, PQ3D improves the state-of-the-art on ScanNet200 by 4.9% (AP25), ScanRefer by 5.4% (acc@0.5), Multi3DRefer by 11.7% (F1@0.5), and Scan2Cap by 13.4% (CIDEr@0.5). Moreover, PQ3D supports flexible inference with individual or combined forms of available 3D representations, e.g., solely voxel input.
Keywords:
3D Vision-Language 3D Scene Understanding Visual Reasoning
@@ 1 Introduction
Recent advancements in embodied artificial intelligence have emphasized the importance of connecting 3D scene understanding with natural language [21, 34, 52, 79, 32]. This step is crucial for embodied agents to understand and execute human instructions in real-world scenarios [5, 59]. In recent years, numerous tasks and datasets for benchmarking 3D scene understanding with languages have been proposed, including 3D semantic segmentation [60], 3D vision-language (3D-VL) reasoning (referring [37, 8, 2, 74, 1], question answering [3, 49, 76], and captioning [12]), and open-vocabulary 3D understanding [27, 64, 39, 25].
@@
We conduct extensive experiments on ten 3D-VL datasets, including ScanNet200 [60] / Replica [63] for instance segmentation, ScanRefer [8] / ReferIt3D [2] / Multi3DRef [74] for visual grounding, ScanQA [3] / SQA3D for question-answering, Scan2Cap [12] for dense captioning, and ObjNav from CortexBench [50] for embodied navigation. The proposed PQ3D achieves impressive results across these tasks, setting new records on most tasks as shown in Fig. 2(a). For example, our model boosts the state-of-the-art on ScanNet200 by 4.9% (AP25), ScanRefer by 5.4% (acc@0.5), Multi3DRef by 11.7% (F1@0.5), and Scan2Cap by 13.4% (CIDEr@0.5). More importantly, PQ3D is the first unified model capable of handling all these tasks simultaneously. The proposed model also shows zero-shot capability with novel prompt types; for instance, we can prompt it with an image sketch to locate all related objects in a scene as in Fig. 1.
Figure 2: Comparison between PQ3D and other models. (a) When comparing PQ3D to other state-of-the-art (SOTA) methods, PQ3D demonstrates superior performance on most tasks. (b) Previous models have been designed for specific tasks and representations, often limiting the potential for developing a unified model. (c) Our PQ3D can flexibly accommodate various input representations, effectively addressing a wide range of tasks.
@@ 3.1 Task Prompt Encoding
There are three widely used representations for 3D scenes: point clouds, voxel grids, and multi-view images, which have their unique advantages in different tasks. For example, voxels offer a uniform, grid-like structure ideal for instance segmentation [60, 62], point clouds provide spatial information crucial for visual grounding [79, 10], and multi-view images include a rich visual context beneficial for open-vocabulary learning [53, 64].
To simultaneously handle various tasks in a 3D scene, we aim to achieve seamless unification of these representations. To accomplish this, we first group the dense 3D points into larger segments through unsupervised graph-based segmentation and masks [23]. The grouped segments are much fewer than the original points, making the training feasible by reducing the number of tokens in cross-attention. Then, we encode these scene representations by the corresponding encoders and pool the features to the segments in total of $M$.
@@ 3.3 Prompt-guided Query Learning
Following [14, 62], we adopt masked attention when cross-attending to the scene features, which restricts the attention to localized features centered around the query for faster convergence and improved performance. To support flexible inference when only some representations are available, we randomly drop out some scene features with rate 0.6 in masked-attention computation during training. After $N$ decoder layers, we expect the final instance queries $\mathbf{Q}$ to have collected enough information for solving the given task.
@@ 4.1 Experimental setting
Training Details The training procedure consists of two stages. In the first stage, we train the model with instance segmentation alone on ScanNet200 for 800 epochs. At this stage, instance segmentation is trained with a classification head on instance queries, instead of prompting object categories. At the second stage, we continue train the model on the full training set using the training objectives in Eq. 7 for 50 epochs. We set the hidden dimension $D$ to 768, and query decoder layer $N$ to 4. We utilize the AdamW optimizer with a learning rate of 1e-4, batch size of 16, $\beta_{1}$ = 0.9, and $\beta_{2}$ = 0.98. The loss balance weights $\lambda_{\mathrm{mask}},\lambda_{\mathrm{gen}}$ are set to 1, and $\lambda_{\mathrm{grd}}$ is set to 10. To further demonstrate the capability of PQ3D, we also transfer it to an embodied agent for object navigation using the ObjNav task from CortexBench [50] and instruction-tune it with a large language model (LLM) Vicuna-7B using the instruction-following dataset from [32]. The whole training process is conducted on four NVIDIA A100 GPUs. More details can be found in the appendix.
@@ 5 Conclusions and Future Works
Acknowledgements We thank the anonymous reviewers for their constructive suggestions. Their insights have greatly improved the quality and clarity of our work. This work was partially supported by the National Science and Technology Major Project (2022ZD0114900). Additionally, Ziyu Zhu and Zhidong Deng were supported by the National Science Foundation of China (NSFC) under Grant No. 62176134 and by research and application on AI technologies for smart mobility funded by SAIC Motor.